OpenStudy (ffasinger):

What is the equation of the following line? Be sure to scroll down first to see all answer options.

OpenStudy (ffasinger):

A y=2x B y-1/2x C y=-4x D y=-1/2 E y=-2x F y=-1/4x

OpenStudy (okdutchman7):

What do you think?

OpenStudy (ffasinger):

no clue

OpenStudy (okdutchman7):

Well you understand that the slope is rise over run, right?

OpenStudy (ffasinger):

is it D!!

OpenStudy (okdutchman7):

That's right. Good Job!